What is the Equation of the line containing (2,-5)and (6,3)

Found 2 solutions by saifoo.khan, stanbon:
Answer by saifoo.khan(162)
Answer by stanbon(75887)

What is the Equation of the line containing (2,-5)and (6,3)
----------------------
slope = (3--5)/(6-2) = 8/4 = 2
-----
Form: y = mx + b
----
Solve for "b":
3 = 2*6 + b
b = -9
-----
Equation:
y = 2x-9
========================
Cheers,
Stan H.
============